# Revisiting Reliability in Large-Scale Machine Learning Research Clusters

### v1

@@ Abstract
We hereby introduce a taxonomy of failures and key reliability metrics, analyze 11 months of data from two state-of-the-art ML environments with over 150 million A100 GPU hours and 4 million jobs. Building on our data, we fit a failure model to project Mean Time to Failure for various GPU scales. We further propose a method to estimate a related metric, Effective Training Time Ratio, as a function of job parameters, and we use this model to gauge the efficacy of potential software mitigations at scale. Our work provides valuable insights and future research directions for improving the reliability of AI supercomputer clusters, emphasizing the need for flexible, workload-agnostic, and reliability-aware infrastructure, system software, and algorithms.
@@ I Introduction
Accelerating innovations towards Artificial General Intelligence is pushing the capability of today’s computing infrastructure, demanding system breakthroughs for model training at-scale. Companies are investing towards large scale clusters with multi-thousand GPUs and fast interconnect networks. For example, Meta introduced its AI supercomputer, interconnecting 2,000 NVIDIA DGX A100 systems (16,000 A100 GPUs) [4] with a 1600 Gb/s InfiniBand network and petabytes of storage capacity [4]. Within just two years, Meta debuted two 24,000-GPU training clusters to accelerate generative AI technology development [1]. At the same time, Google invested in the next generation of Tensor Processing Units (TPUs) that form the foundation of its AI supercomputers [27, 26]. Each TPU v5p pod constitutes 8,960 chips, which communicate via the inter-chip interconnect of 4,800 Gb/s/chip in a 3D torus topology [2]. Such heavy infrastructure investments inevitably stress the limits of the existing systems stack.
@@
With the rise of Large Language Models (LLMs)—Megascale [25], LLaMa [41], Gemini [17], GPT4 [35]—ML training shifted the scale of a single training job from tens to tens of thousands of accelerators, presenting concrete use-cases for the aforementioned investments. At such scale, failures are not a matter of if, but a matter of when. Thus, system design entails new and unexpected challenges in the operation, efficiency and reliability of a cluster, creating the need for new solutions. Indeed, hardware failures [17, 33] are just some of the issues that are individually rare yet become increasingly likely at scale, involving solutions that span from hardware to the design of a machine learning cluster scheduler.
@@
In this paper, we present our infrastructure experience toward training a plethora of large scale models, including earlier foundation models [41] as their usage became prevalent, with largest jobs utilizing 4k GPUs or more. Unlike prior work, our hardware and software infrastructure is tailor-designed to be capable of serving a diverse set of workloads—4k GPU jobs constitute less than 1% of our jobs while consuming 12% of the GPU resources at the cluster level. Our experience catering to both large- and small-scale jobs demonstrates diversity in infrastructure needs that is rarely observed in more specialized clusters devoted to LLMs only.
@@
Understanding the underlying causes of job failures—let it be hardware, system software, applications or some combinations of the above—is key to improve training reliability, and advance large model development. In this paper, we present 11 months of data collected from state-of-the-art AI researcher clusters with >80% utilization. The results based on real-world training systems highlight the diversity of research workloads across 2 clusters, spanning 24k of NVIDIA A100 GPUs. Our primary focus is on job-level failures. We primarily view failures through the lens of the scheduler and server-level health checks. We additionally provide some network-level reliability experience. Finally, we share lessons we have learned in mitigating failures at scale, tracking reliability metrics, making infrastructure changes, and diagnosing common application pitfalls, culminating in suggestions for future opportunities. In doing so, we provide and analyze server-level component failure rates, including Mean Time to Failure (MTTF) projections.
@@
    Pinpointing reliability improvement opportunities based on an analysis of deployed ML training systems. The 11-month data from training various ML jobs in two state-of-the-art machine learning environments spans over 150 million A100 GPU hours and 4 million jobs. We find that jobs in our research cluster are more diverse than implied by LLM workloads, motivating workload agnostic infrastructure techniques.
@@ II System Infrastructure
In this section, we describe how workloads influence the design of our clusters. While clusters can be specialized to optimize toward a specific workload, research clusters are, by definition, expected to have constantly changing workloads with potentially unforeseen needs. Therefore, we believe research clusters should be general, maximize productivity, and minimize incidental complexity. Our two sister clusters, RSC-1 and RSC-2, follow the same design template discussed below. RSC-1 is a general ML cluster (e.g., training some of the prominent LLMs) of 16k GPU size, while RSC-2 focuses on vision applications and is of 8k GPU size. As we discuss later (§III), the workload differences manifest in different usages—for example, workloads on RSC-2 have a significant tilt towards 1-GPU jobs, along with jobs going up to 1k GPU size.
@@ II-A Scheduler and Storage Infrastructure Overview
Our design of RSC-1 and RSC-2 prioritized ease of use while favoring simplicity. The benefits of our design is that the entire stack is mature and does not require extensive custom datacenter designs, reducing our time-to-market to only 1.5 years. Additionally, we aim to provide users with the requested number of GPUs with no strings attached to maximize productivity—users do not have to deal with complexity in the form of novel hardware or virtualization. Figure 1 provides an overview of how users interact with both our clusters. Users submit a job, which is comprised of many tasks, each of which can run on node’s GPU.
Scheduler. Leaning into the High-Performance Computing (HPC) stack, our clusters use the Slurm [45] scheduler on top of bare-metal allocations. The cluster is configured such that groups of users have a maximum quota of GPUs that is determined by a project-specific allocation. Users submit jobs using shell scripts (sbatch) or Python wrappers (submitit [12]). Slurm, in turn, attempts to co-locate the tasks given the physical network topology. Jobs are eligible to be preempted after running for 2 hours, and they have a maximum lifetime of 7 days. The scheduler attempts to schedule jobs based on priority order, which is a function of many variables, including the project’s allocation and the job’s age [5].
@@ II-C Observations on Cluster Infrastructure
Observation 1: Cluster uptime is critical. Our clusters are fully loaded. Any downtime results in excessive queuing and is considered a major event. The cluster must adapt to failures online, and ideally auto-requeue infrastructure-related failures.
@@
Slurm can run checks before and after a job runs [10]. Moreover, we have health checks that are periodically scheduled to run every five minutes, and return codes indicating success, failure, or warning. Each health check examines some aspect of node health, spanning from GPU errors (e.g. XID errors [9]) to file system mounts, and services status (i.e., scheduler). Note that checks can have overlapping signals into a failure domain. For example, a PCIe failure indicates that the GPU is inaccessible, even if the GPU did not incur the corresponding XID event itself. This situation occurs in our logs 57% of the time on RSC-1 (37% on RSC-2). Therefore, even if one check does not fire when it should, another overlapping check would hopefully catch the failure. The most extreme case of this is NODE_FAIL, which acts as a catch-all via Slurm heartbeats when a node becomes unresponsive to other health checks that are running on the node itself.
Periodic health checks are essential to prevent repeated job failures from same unhealthy nodes. Additionally, the checks are tuned to have a low false positive rate—they have been previously calibrated such that less than 1% of successfully completed jobs observe a failed health check, note though that we can only observe correlations and not causations.
Different checks have different severity. High severity check failures will immediately signal a scheduler handler to remove the node and reschedule all jobs executing on the node, while lower severity checks will signal to the scheduler to remove the node for remediation after jobs running on the node have finished, successfully or unsuccessfuly. In the first category of check are the following: GPU not accessible, an NVLink error, uncorrectable ECC, failed row-remaps, PCI or IB link errors or block device errors, and missing mountpoints. Nodes that are not failing any health checks are available for scheduling jobs. When a health check fails for a node, the node will be transitioned to a remediation state and will be unavailable for scheduling until it is fixed and all checks are passing. The transition to remediation can happen either immediately (for high severity checks) or after the current job finishes.
@@ II-D Metrics
There are three critical metrics to understanding how an ML cluster is performing: Effective Training Time Ratio (ETTR), Model Flops Utilization (MFU), and Goodput.
Effective Training Time Ratio (ETTR). ETTR is defined as the ratio of productive runtime to the available wallclock time of a job run. A job run consists of one or more scheduler jobs related to the same logical job [37]. For example, a multi-week LLM pretraining run may consist of multiple different jobs demarcated by pre-emptions and infrastructure failures (we ignore impact to ETTR from userspace failures to focus on impact from cluster stability only). The available wallclock time of a job run is defined as the total time a job in the multi-job run was either 1 scheduled or 2 eligible to be scheduled but waiting in the queue. Productive runtime refers to scheduled time during which meaningful progress is being made for the workload. The exact definition of productive runtime is open to interpretation depending on context, but we consider two sources of unproductive scheduled time:
@@
Both of these are highly job dependent, and we currently lack a reliable way for tracking either at scale with confidence. However, we treat these as free parameters to explore, filling in with reasonable values we have encountered anecdotally in collaborating with various research teams.
@@
ETTR-like metrics, such as tracking job runtime until failure, were initially used for tracking the training efficiency of our LLMs, like LLaMa [41], as infrastructure issues were iteratively diagnosed. Since then, such metrics were generalized to ETTR and continue to be useful for more recent LLMs [33] outside of the presented clusters. For instance, Google Cloud defines an equivalent metric that they term “Runtime Goodput” [3]. To differentiate between the goodput metric we refer to in this paper (which only includes impact from wasted compute and not from e.g. wait time), we coin the term Effective Training Time Ratio (ETTR).
Model Flops Utilization (MFU). A related metric used throughout the industry is model flops utilization (MFU) [29]. This metric corresponds to the number of FLOPs a model theoretically utilizes compared to hardware peak FLOPs. Since ETTR is model agnostic and considers job restarts and queueing, it presents a more general measure of cluster reliability and goodput achieved. However, unlike MFU, ETTR does not capture degraded performance or suboptimal implementations. Both are comparable in the sense that they measure the ratio of overhead runtime to ideal and are between 00 and $1$, though ETTR is typically higher (e.g., over 80%, while LLM MFU is smaller, for example 38-43% for LLaMa 3 [33]).
Goodput. Both of the above metrics can be viewed as per-job efficiency metrics. The cluster as a whole can be measured in terms of goodput, which is the amount of productive work completed in aggregate per unit time. The goodput can be normalized by the maximum possible goodput to produce a utilization in the range 0 to 1. The clusters discussed in this paper operate at high utilization (so potential goodput is only limited by capacity rather than available work), and thus job preemption, resource fragmentation, and failures are the dominant sources of lost goodput.
@@ II-E Failure Taxonomy
Failure attribution is the process of assigning blame for a job failure to a cause. Our experience indicates that failure attribution is a challenging and noisy process. For instance, NCCL timeouts are a relatively common occurrence [21]. In PyTorch, a NCCL timeout occurs whenever a rank observes that a collective operation, such as an AllReduce, has not completed within a several minutes. While this can mean that some network issue occurred, it can also mean that some other rank simply never started that same operation because it was, for example, stuck trying to load data for the next iteration. In this case, the rank that times out is fully functional. The rank at fault may itself be unresponsive either due to a user software or due to infrastructure error (which itself can occur at a link or switch level). Tracing the root cause from user-level stack traces would require potentially many layers of precise and distributed logging, spanning from the ML application down to distributed collectives and low-level infrastructure.
@@ III Understanding Lay of the Land of Large-Scale ML Training Clusters
Our analysis is based on two research clusters and spans 11 months of measurement data. It builds on the terminology of the Slurm scheduler and the node-level health-checks discussed previously (§II-A). Note that the clusters discussed in this section are over-provisioned, so project-level QoS and allocations are major factors in determining which jobs run.
@@
Scheduler Job Status Breakdown: A Slurm job can be CANCELLED, COMPLETED, OUT_OF_MEMORY, or FAILED because the application returned a non-zero exit code, NODE_FAIL because of a faulty node, PREEMPTED in favor of a higher priority job, REQUEUED, or TIMEOUT. Figure 3 illustrates the scheduler job status breakdown for the RSC-1 cluster. 60% of scheduled jobs completed. 24% and 0.1% of jobs failed because of FAILED and NODE_FAIL, respectively. 10% of jobs were pre-empted, 2% requeued, 0.1% ran out of memory, and 0.6% timed-out.
@@
Job-Level Failure Characterization: Attributed hardware failures can be broken down by attributed cause. These causes can be further subdivided by server-level components, such as the GPU, the network, and various system components, such as the filesystem. We show such GPU-hour normalized failure rates for RSC-1and RSC-2 in Figure 4. We attribute a failure to a cause if the cause was detected within the last 10 minutes or 5 minutes after a failing jobs lifetime (FAILED or NODE_FAIL). Note that we report the most likely cause of failure according to heuristics we developed indicating whether a node should be isolated for remediation. Some failures have multiple attributions (see Figure 4). Some NODE_FAIL events are not associated with any health checks (c.f., [30]), likely because the node itself became unresponsive. IB Links, filesystem mounts, GPU memory errors, and PCIe errors contribute heavily to the failure rates, however for IB Links in particular this seems to be dominated by a short period of many IB Link related job failures from a handful of nodes in the summer of 2024 as shown in Figure 5. GSP Timeouts were caused by a code regression that was fixed with a driver patch (see Figure 5).
@@
Evolution of failure rate over time: We now turn our analysis to larger jobs, therefore switching to node-level (rather than GPU-level) analysis. In Figure 5, we illustrate how failure rate looks for RSC-1 over the last year (showing a 30 day rolling average), illustrating:
  • Failure rate is constantly changing. We see periods where e.g., failure rate is $\sim$2.5 failures per node day on RSC-1 and periods where failure rate spikes as high as $\sim$17.5 failures per node day (an order of magnitude higher).
  • Failure modes ebb and flow. In late 2023, XID errors from a driver bug were the dominant source of job failures on RSC-1; this issue was resolved. In spring of 2023, after adding a new health check for mounts that were downing nodes, this became a key failure mode on RSC-1. In early summer of 2024, a spike of IB Link failures on a small number of offending nodes temporarily drove up the failure rate on both clusters.
  • New health checks expose new failure modes. We mark the time that new health checks were added to the cluster. The addition of a new health check, usually in response to an anecdotal report of a previously unchecked failure mode, has a tendency to cause an apparent increase in failure rate simply because we suddenly are able to see a failure mode that was likely previously present.
@@
Training Job Diversity: We have a diverse collection of training jobs in sizes and in the overall GPU hours. The scheduler must consider job size diversity and the corresponding training time to balance between training time performance, fairness of individual training jobs, and overall cluster utilization.
Figure 6 depicts the distribution of job size for the RSC-1 cluster. More than 40% of training jobs use a single GPU for development or for model evaluation. There are only a few large-scale training jobs, utilizing thousand GPUs in the research clusters. In the same figure, we also illustrate the corresponding GPU time normalized to the 1-GPU training job. Despite the many 1-GPU training jobs, more than 66% and 52% of the overall GPU time comes from 256+ GPU jobs for the RSC-1 and RSC-2 clusters, respectively. Compared with production model training for LLaMa 3.0 [33], the diversity of training job sizes and the respective training time in the research clusters poses unique challenges to the design of an effective ML scheduler.
@@
MTTF Decreases at Scale. Figure 7 illustrates that the mean-time-to-failure (MTTF) of 1024-GPU jobs is 7.9 hours—roughly 2 orders-of-magnitude lower than 8-GPU jobs (47.7 days). As shown in §II-E, training failures stem from various factors. From user programs to system software and hardware faults. Empirically, hardware reliability shrinks inversely proportional to number of GPUs, with more consistent trends starting at 32 GPUs. 90% confidence intervals generated by fitting a Gamma distribution.
@@
Figure 7: MTTF analysis by job sizes for RSC-1 and RSC-2, rounded up to the next multiple of 8 GPUs. CI: Confidence Interval. MTTF decreases predictably with scale.

### v2

@@ Abstract
We hereby introduce a taxonomy of failures and key reliability metrics, analyze 11 months of data from two state-of-the-art ML environments with 4 million jobs and over 150 million A100 GPU hours. Building on our data, we fit a failure model to project Mean Time to Failure for various GPU scales. We further propose a method to estimate a related metric, Effective Training Time Ratio, as a function of job parameters, and we use this model to gauge the efficacy of potential software mitigations at scale. Our work provides valuable insights and future research directions for improving the reliability of AI supercomputer clusters, emphasizing the need for flexible, workload-agnostic, and reliability-aware infrastructure, system software, and algorithms.
@@ I Introduction
Accelerating innovations towards Artificial General Intelligence is pushing the capability of today’s computing infrastructure, demanding system breakthroughs for model training at-scale. Companies are investing in large-scale clusters with thousands of GPUs and fast interconnect networks. For example, Meta introduced its AI supercomputer, interconnecting 2,000 NVIDIA DGX A100 systems (16,000 A100 GPUs) [4] with a 1600 Gb/s InfiniBand network and petabytes of storage capacity [4]. Within just two years, Meta debuted two 24,000-GPU training clusters to accelerate generative AI technology development [1]. At the same time, Google invested in the next generation of Tensor Processing Units (TPUs) that form the foundation of its AI supercomputers [40, 39]. Each TPU v5p pod constitutes 8,960 chips, which communicate via the inter-chip interconnect of 4,800 Gb/s/chip in a 3D torus topology [2]. Such heavy infrastructure investments inevitably stress the limits of the existing systems stack.
@@
With the rise of Large Language Models (LLMs)—Megascale [38], LLaMa [56], Gemini [28], GPT4 [49]—ML training shifted the scale of a single training job from tens to tens of thousands of accelerators, presenting concrete use-cases for the aforementioned investments. At such scale, failures are not a matter of if, but a matter of when. Thus, system design entails new and unexpected challenges in the operation, efficiency, and reliability of a cluster, creating the need for new solutions. Indeed, hardware failures [28, 47] are just some of the issues that are individually rare yet become increasingly likely at scale, involving solutions that span from hardware to the design of a machine learning cluster scheduler.
@@
In this paper, we present our infrastructure experience toward training a plethora of large-scale models, including earlier foundation models [56] as their usage became prevalent, with the largest jobs utilizing 4k GPUs or more. Unlike prior work, our hardware and software infrastructure is tailor-designed to be capable of serving a diverse set of workloads—4k GPU jobs constitute less than 1% of our jobs while consuming 12% of the GPU resources at the cluster level. Our experience catering to both large- and small-scale jobs demonstrates diversity in infrastructure needs that is rarely observed in more specialized clusters devoted to only LLMs.
@@
Understanding the underlying causes of job failures—let it be hardware, system software, applications, or some combinations of the above—is key to improving training reliability and advancing large model development. In this paper, we present 11 months of data collected from state-of-the-art AI research clusters with >80% utilization. The results based on real-world training systems highlight the diversity of research workloads across 2 clusters, spanning 24k of NVIDIA A100 GPUs. Our primary focus is on job-level failures. We primarily view failures through the lens of the scheduler and server-level health checks. We additionally provide some network-level reliability experience. Finally, we share lessons we have learned in mitigating failures at scale, tracking reliability metrics, making infrastructure changes, and diagnosing common application pitfalls, culminating in suggestions for future opportunities. In doing so, we provide and analyze server-level component failure rates, including Mean Time to Failure (MTTF) projections.
@@
    Pinpointing reliability improvement opportunities based on an analysis of deployed ML training systems. The 11-month data from training various ML jobs in two state-of-the-art machine learning environments spans 4 million jobs and over 150 million A100 GPU hours. We find that jobs in our research cluster are more diverse than implied by LLM workloads, motivating workload agnostic infrastructure techniques.
@@ II System Infrastructure
In this section, we describe how workloads influence the design of our clusters. While clusters can be specialized to optimize toward a specific workload, research clusters are, by definition, expected to have constantly changing workloads with potentially unforeseen needs. Therefore, we believe that research clusters should be general, maximize productivity, and minimize incidental complexity. Our two sister clusters, RSC-1 and RSC-2, follow the same design template discussed below. RSC-1 is a general ML cluster (e.g., training some of the prominent LLMs) of 16k GPU size, while RSC-2 focuses on vision applications and is of 8k GPU size. As we discuss later (§III), the workload differences manifest in different usages—for example, workloads on RSC-2 have a significant tilt towards 1-GPU jobs, along with jobs going up to 1k GPU size.
@@ II-A Scheduler and Storage Infrastructure Overview
Our design of RSC-1 and RSC-2 prioritized ease of use while favoring simplicity. The benefit of our design is that the entire stack is mature and does not require extensive custom datacenter designs, reducing our time-to-market to only 1.5 years. Additionally, we aim to provide users with the requested number of GPUs with no strings attached to maximize productivity—users do not have to deal with complexity in the form of novel hardware or virtualization. Figure 1 provides an overview of how users interact with both our clusters. Users submit a job, which is comprised of many tasks, each of which can run on the GPUs of a node.
Scheduler. Leaning into the High-Performance Computing (HPC) stack, our clusters use the Slurm [62] scheduler on top of bare-metal allocations. The cluster is configured such that groups of users have a maximum quota of GPUs that is determined by a project-specific allocation. Users submit jobs using shell scripts (sbatch) or Python wrappers (submitit [13]). Slurm, in turn, attempts to co-locate the tasks given the physical network topology. Jobs are eligible to be preempted after running for two hours, and they have a maximum lifetime of seven days. Slurm attempts to schedule jobs based on priority order, which is a function of many variables, including the project’s allocation and the job’s age [5].
@@ II-C Observations on Cluster Infrastructure
Observation 1: Cluster uptime is critical. Our clusters are fully loaded. Any downtime results in excessive queueing and is considered a major event. The cluster must adapt to failures online and ideally auto-requeue infrastructure-related failures.
@@
Slurm can run checks before and after a job runs [10]. Moreover, we have health checks that are periodically scheduled to run every five minutes and return codes indicating success, failure, or warning. Each health check examines some aspect of node health, spanning from GPU errors (e.g. XID errors [9]) to file system mounts, and services status (i.e., scheduler). Note that checks can have overlapping signals into a failure domain. For example, a PCIe failure indicates that the GPU is inaccessible, even if the GPU did not incur the corresponding XID event itself. This situation occurs in our logs 57% of the time on RSC-1 (37% on RSC-2). Therefore, even if one check does not fire when it should, another overlapping check would hopefully catch the failure. The most extreme case of this is NODE_FAIL, which acts as a catch-all via Slurm heartbeats when a node becomes unresponsive to other health checks that are running on the node itself.
Periodic health checks are essential to prevent repeated job failures from the same unhealthy nodes. The checks are tuned to have a low false positive rate—they have been previously calibrated such that less than 1% of successfully completed jobs observe a failed health check, though note that we can only observe correlations and not causations.
Different checks have different severity. High severity check failures will immediately signal a scheduler handler to remove the node and reschedule all jobs executing on the node, while lower severity checks will signal to the scheduler to remove the node for remediation after jobs running on the node have finished, successfully or unsuccessfuly. In the first category of check are the following: GPU not accessible, an NVLink error, uncorrectable ECC, failed row-remaps, PCI or IB link errors, block device errors, and missing mountpoints. Nodes that are not failing any health checks are available for scheduling jobs. When a health check fails for a node, the node will transition to a remediation state and will become unavailable for scheduling until it is fixed and all checks are passing. The transition to remediation can happen either immediately (for high severity checks) or after the current job finishes.
@@ II-D Metrics
There are three critical metrics we consider in this paper for understanding how an ML cluster is performing: Effective Training Time Ratio (ETTR), Goodput, and Mean Time to Failure (MTTF).
Effective Training Time Ratio (ETTR). ETTR is defined as the ratio of productive runtime to available wallclock time of a job run. A job run consists of one or more scheduler jobs related to the same logical job [52]. For example, a multi-week LLM pretraining run may consist of multiple different jobs demarcated by pre-emptions and infrastructure failures (ETTR attempts to ignore the impact from userspace failures to focus on impact from cluster stability only). The available wallclock time of a job run is defined as the total time a job in the multi-job run was either 1 scheduled or 2 eligible to be scheduled but waiting in the queue. Productive runtime refers to scheduled time during which meaningful progress is being made for the workload. The exact definition of productive runtime is open to interpretation depending on context, but we consider three sources of unproductive scheduled time:
@@
  3. 3.
    Checkpoint overhead: The time checkpointing adds to job runtime.
All of these are highly job dependent, and we currently lack a reliable way for tracking either at scale with confidence. However, we treat these as free parameters to explore, filling in with reasonable values we have encountered anecdotally in collaborating with various research teams.
@@
ETTR-like metrics, such as tracking job runtime until failure, were initially used for tracking the training efficiency of our LLMs, like LLaMa [56], as infrastructure issues were iteratively diagnosed. Since then, such metrics were generalized to ETTR and continue to be useful for more recent LLMs [47] outside of the presented clusters. For instance, Google Cloud defines a similar metric that they term “Runtime Goodput” [3]. To differentiate between the goodput metric we refer to in this paper (which only includes impact from wasted compute and not from e.g. wait time), we use the term Effective Training Time Ratio (ETTR). Similarly, note that the ETTR we use differs from other definitions [61] in that we model the wait time found in multi-tenant clusters.
Other potential metrics for characterizing model performance include Model Flops Utilization (MFU) [42, 21], which we leave out of this paper. MFU corresponds to the number of FLOPs a model theoretically utilizes compared to the hardware peak FLOPs, making it difficult to apply generally across an entire cluster.
Goodput. ETTR and MFU can be viewed as per-job efficiency metrics. The cluster as a whole can be measured in terms of goodput, which is the amount of productive work completed in aggregate per unit time. The goodput can be normalized by the maximum possible goodput to produce a utilization in the range 0 to 1. The clusters discussed in this paper operate at high utilization (so potential goodput is limited more by capacity rather than available work), and thus job preemption, resource fragmentation, and failures are the dominant sources of lost goodput. While we use goodput to communicate loss in certain restricted scenarios in this paper, we focus on ETTR as the main measure of job productivity on our clusters.
Mean Time to Failure (MTTF). A key statistic in any reliability study is the Mean Time to Failure (MTTF), a measure of how often failures occur. It is the amount of measured system time divided by the amount of failures. The failure rate is the inverse. The MTTF ranges from 0 to $\infty$ and gets smaller as sources of failure sum to higher total failure rates and thus lower MTTF. MTTF can be used to configure the optimal checkpoint strategy under nonzero checkpoint overhead [23, 63].
@@ II-E Failure Taxonomy
Failure attribution is the process of assigning blame for a job failure to a cause. Our experience indicates that failure attribution is a challenging and noisy process. For instance, NCCL timeouts are a relatively common occurrence [32]. In PyTorch, a NCCL timeout occurs whenever a rank observes that a collective operation, such as an All-Reduce, has not completed within several minutes. While this can mean that some network issue occurred, it can also mean that some other rank simply never started that same operation because it was, for example, stuck trying to load data for the next iteration. In this case, the rank that times out is fully functional. The rank at fault may itself be unresponsive either due to a user software or due to infrastructure error (which itself can occur at a link or switch level). Tracing the root cause from user-level stack traces would require potentially many layers of precise and distributed logging, spanning from the ML application down to distributed collectives and low-level infrastructure.
@@ III Understanding Lay of the Land of Large-Scale ML Training Clusters
Our analysis is based on two research clusters and spans 11 months of measurement data. It builds on the terminology of the Slurm scheduler and the node-level health-checks discussed previously (§II-A). Note that the clusters discussed in this section are over-provisioned, and project-level QoS and allocations are major factors in determining which jobs run.
@@
Scheduler Job Status Breakdown. A Slurm job can be CANCELLED, COMPLETED, OUT_OF_MEMORY, FAILED because the application returned a non-zero exit code, NODE_FAIL because of a faulty node, PREEMPTED in favor of a higher priority job, REQUEUED, or TIMEOUT. Figure 3 illustrates the scheduler job status breakdown for the RSC-1 cluster. 60% of scheduled jobs completed. 24% and 0.1% of jobs failed because of FAILED and NODE_FAIL, respectively. 10% of jobs were pre-empted, 2% requeued, 0.1% ran out of memory, and 0.6% timed-out.
@@
Job-Level Failure Characterization. Attributed hardware failures can be broken down by attributed cause. These causes can be further subdivided by server-level components, such as the GPU, the network, and various system components, such as the filesystem. We show such GPU-hour normalized failure rates for RSC-1 and RSC-2 in Figure 4. We attribute a failure to a cause if the cause was detected within the last 10 minutes or 5 minutes after a failing jobs lifetime (FAILED or NODE_FAIL). Note that we report the most likely cause of failure according to heuristics we developed indicating whether a node should be isolated for remediation. Some failures have multiple attributions (see Figure 4). Some NODE_FAIL events are not associated with any health checks (c.f., [43]), likely because the node itself became unresponsive. IB Links, filesystem mounts, GPU memory errors, and PCIe errors contribute heavily to the failure rates, however for IB Links in particular this seems to be dominated by a short period of many IB Link related job failures from a handful of nodes in the summer of 2024 as shown in Figure 5. GSP Timeouts were caused by a code regression that was fixed with a driver patch (see Figure 5).
@@
Evolution of Failure Rate over Time. We now turn our analysis to larger jobs, therefore switching to node-level (rather than GPU-level) analysis. In Figure 5, we show how failures manifest for RSC-1 over the last year (plotting failure rates using a 30 day rolling average), illustrating:
  • Failure rate is constantly changing. We see periods where e.g., failure rate is $\sim$2.5 failures per 1000 node-days on RSC-1 and periods where failure rate spikes as high as $\sim$17.5 failures per 1000 node-days (an order of magnitude higher).
  • Failure modes ebb and flow. In late 2023, XID errors from a driver bug were the dominant source of job failures on RSC-1; this issue was resolved. In spring of 2024, after adding a new health check for mounts that were downing nodes, this became a key failure mode on RSC-1. In early summer of 2024, a spike of IB Link failures on a small number of offending nodes temporarily drove up the failure rate on both clusters.
  • New health checks expose new failure modes. We mark the time that new health checks were added to the cluster. The addition of a new health check, usually in response to an anecdotal report of a previously unchecked failure mode, has a tendency to cause an apparent increase in failure rate, simply because we are suddenly able to see a failure mode that was likely previously present.
@@
Training Job Diversity. We have a diverse collection of training jobs in terms of job size and the overall consumed GPU hours. The scheduler must consider job size diversity and the corresponding training time to balance between training time performance, fairness of individual training jobs, and overall cluster utilization.
Figure 6 depicts the distribution of job size for the RSC-1 cluster. More than 40% of training jobs use a single GPU for development or for model evaluation. There are only a few large-scale training jobs, which utilize thousands of GPUs in the research clusters. In the same figure, we also illustrate the corresponding percentage of GPU time consumed by the jobs. Despite many 1-GPU training jobs, more than 66% and 52% of the overall GPU time comes from 256+ GPU jobs for the RSC-1 and RSC-2 clusters, respectively. Compared with production model training for LLaMa 3.0 [47], the diversity of training job sizes and the respective training time in the research clusters poses unique challenges to the design of an effective ML scheduler.
@@
MTTF Decreases at Scale. Figure 7 illustrates that the mean-time-to-failure (MTTF) of 1024-GPU jobs is 7.9 hours—roughly 2 orders-of-magnitude lower than 8-GPU jobs (47.7 days). As shown in §II-E, training failures stem from various factors, ranging from user programs to system software to hardware faults. Empirically, hardware reliability shrinks inversely proportional to the number of GPUs, with more consistent trends starting at 32 GPUs. 90% confidence intervals are generated by fitting a Gamma distribution.
@@
Figure 7: MTTF analysis by job sizes for RSC-1 and RSC-2, rounded up to the next multiple of 8 GPUs. CI: Confidence Interval (90%). MTTF decreases predictably with scale.